OpenStudy (ffasinger):

What is the value of x in the rhombus below? A. 47 B. 43 C. 21 D. 25

OpenStudy (soprano.h.d0816):

First, you would set them equal to each other

OpenStudy (ffasinger):

how do you do that @soprano.h.d0816

OpenStudy (soprano.h.d0816):

2x+1 = 2x+5

OpenStudy (ffasinger):

k did that

OpenStudy (soprano.h.d0816):

Holy gosh! I so read that wrong. Scratch what I said!! Lol

OpenStudy (ffasinger):

lol okay

OpenStudy (soprano.h.d0816):

What you would actually do is: 2x+1+2x+5=180 and combine like signs

OpenStudy (ffasinger):

so it would be 4x +6=180

OpenStudy (soprano.h.d0816):

Then subtract 6 from both sides

OpenStudy (ffasinger):

174

OpenStudy (ffasinger):

43.5

OpenStudy (soprano.h.d0816):

So: 4x = 174 Divide both sides by 4 and x = ??

OpenStudy (soprano.h.d0816):

Good job!!

OpenStudy (ffasinger):

it was wrong but ill give you a medal anyways for helping.

OpenStudy (soprano.h.d0816):

Oh. Dang... but thanks(: